Mean proportional of a and b is 14. If a+b=53, then find a and b.
Answer by stanbon(75887)

Mean proportional of a and b is 14. If a+b=53, then find a and b.
---
Equations:
a/14 = 14/b
-----
a+b = 53
==============
a = b
----
2a = 53
a = 26.5
b = 26.5
==============
Cheers,
Stan H.